40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 9 11:33:49 2017, gautam, Update, Electronics, MC2 de-whitening  MC2deWhites.pdfMC2Coils.pngMC2stab.pdf
    Reply  Wed Aug 9 12:05:57 2017, rana, Update, Electronics, data archiving 
Message ID: 13174     Entry time: Wed Aug 9 11:33:49 2017     Reply to this: 13176
Author: gautam 
Type: Update 
Category: Electronics 
Subject: MC2 de-whitening  

Summary:

The analog de-whitening filters for MC2 are different from those on the other optics (i.e. ITMs and ETMs). They have one complex pole pair @7Hz, Q~sqrt(2), one complex zero pair @50Hz, Q~sqrt(2), one real pole at 2.5kHz, and one real zero @250Hz (with a DC gain of 10dB).

Details:

I took the opportunity last night to measure all 4 de-whitening channel TFs. Measurements and overlaid LISO fits are seen in Attachment #1. 

The motivation behind this investigation was that last week, I was unable to lock the IMC to one of the arms. In the past, this has been done simply by routing the control signal of the appropriate arm filter bank (e.g. C1:LSC-YARM_OUT) to MC2 instead of ETMY via the LSC output matrix (if the matrix element to ETMY is 1, the matrix element to MC2 is -1).

Looking at the coil output filter banks on the MC2 suspension MEDM screen (see Attachment #2), the positions of filters in the filter banks is different from that on the other optics. In general, the BIO outputs of the DAC are wired such that disengaging FM9 on the MEDM screen engages the analog de-whitening path. FM10 then has the inverse of the de-whitening filter, such that the overall TF from DAC to optic is unity. But on MC2, these filters occupy FM7 and FM8, and FM9 was originally a 28Hz Elliptic Low-pass filter.

So presumably, I was unable to lock the IMC to an arm because for either configuration of FM9 (ON or OFF), the signal to the optic was being aggressively low-passed. To test this hypothesis, I simply copied the 28Hz elliptic to FM6, put a gain of 1 on FM9, left it engaged (so that the analog path TF is just flat with gain x3), and tried locking the IMC to the arm again - I was successful. See Attachment #3 for comparison of the control signal spectra of the X-arm control signal, with the IMC locked to the Y-arm cavity.

In this test, I also confirmed that toggling FM9 in the coil output filter banks actually switches the analog path on the de-whitening boards.

Since I now have the measurements for individual channels, I am going to re-configure the filter arrangement on MC2 to mirror that on the other optics. 


Unrelated to this work: the de-whitening boards used for MC1 and MC3 are D000316, as opposed to D000183 used for all other SOS optics. From the D000316 schematic, it looks like the signals from the AI board are routed to this board via the backplane. I will try squishing this backplane connector in the hope it helps with the glitching MC1 suspension.


GV Aug 13 11:45pm - I've made a DCC page for the MC2 dewhitening board. For now, it has the data from this measurement, but if/when we modify the filter shape, we can keep track of it on this page (for MC2 - for the other suspensions, there are other pages). 

Attachment 1: MC2deWhites.pdf  153 kB  | Hide | Hide all
MC2deWhites.pdf
Attachment 2: MC2Coils.png  30 kB  Uploaded Wed Aug 9 12:47:55 2017  | Hide | Hide all
MC2Coils.png
Attachment 3: MC2stab.pdf  40 kB  Uploaded Wed Aug 9 12:48:11 2017  | Hide | Hide all
MC2stab.pdf
ELOG V3.1.3-